OpenStudy (anonymous):

Find a polynomial for the perimeter and for the area one side is d+5 the other side is d

OpenStudy (anonymous):

this is for a triangle that I am looking for

OpenStudy (anonymous):

What kind of triangle?

OpenStudy (anonymous):

I am sorry it is for a rectangle

OpenStudy (anonymous):

Perimeter is just the sum of the sides. so 2(d + 5) + 2(d) = p so 4d + 10 is the perimeter. Area you multiply both sides. (d+5)(d) = A d^2 + 5d = A